OpenStudy (anonymous):

(6x+7)(3x+2) how to solve this one with the steps I have the answer already .

OpenStudy (angela210793):

18x^2+12x+21x+14=18x^2+33x+14 D=b^2-4ac=33^2-4*18*14 If D>0 then X1=(-b+sqrD)/2a X2=(-b-sqrD)/2a if D<0 no solution If D=0 X1=X2=-b/2a